OpenStudy (anonymous):

If BX = 8, BZ = 3, DC = 5, then XC =

OpenStudy (anonymous):

in XC = ?, you need to express C in terms of X or x in terms of C you can do this by working your way through the other equations. From the first equation, you know that b = 8/x. Substitute that into the second equation bd=3, and you get 8d/x=3 now solve for d (multiply both sides by x/8) and you get d=3x/8 Substitute that into the third equation you get 3cx/8 = 5 Solve for cx and you get cx=40/3